OpenStudy (itrymath):

Select the equation of a line that is perpendicular to the line on the graph and passes through the point (-1, 2). (1 point) y = - 1/2 x + 2 y = - 2x + 2 y = 2x + 4 y = 2x - 1

OpenStudy (itrymath):

@Will.H

OpenStudy (will.h):

so 1st we need to find the slope of the line in the graph (-2,0) (0,-1) Slope = -1 - 0/0 - (-2) Slope = -1/2 Now since they said "perpendicular to the line in the graph" that means the slope of the other line will be a number that when you multiply to the 1st slope the answer will be -1 Normally you just have to flip the slope of the 1st line since the slope of the 1st line is -1/2 then the other slope will be 2 since it has to be perpendicular proof: -1/2 * 2 = -1 so indeed the slope of the other line is 2 and" it passes through the point (-1,2)" so now we use the form below y - y1 = m(x - x1) y - 2 = 2(x - (-1)) y - 2 = 2(x + 1) y - 2 = 2x + 2 y = 2x + 4 hope that helps

OpenStudy (itrymath):

it does , im just going to review it because its so much

OpenStudy (will.h):

let me make it easier parallel lines : same slope perpendicular lines: slope with a multiplication process resulting -1 Example if we have a slope of -2/3 the other slope will be 3/2 (you flip the fraction, in other words change the sign and change the places of numerator and denominator)

OpenStudy (itrymath):

wait how do you know the other slope will be 2?

OpenStudy (itrymath):

@Will.H

OpenStudy (will.h):

The lines are perpendicular so therefor when you multiply them the result must be -1 right? make that as equation (-1/2)x = -1 where x is the other slope solve for x x = -1/-0.5 x = 2

OpenStudy (will.h):

OR basically as i mentioned before you can flip the fraction of the 1st slope the 1st slope is -1/2 flip that (switch the numerator and denominator and change the sign from negative to positive) therefor it is 2

OpenStudy (will.h):

Remember you can do that only if the lines are perpendicular

OpenStudy (will.h):

tag me if you need any additional help